OpenStudy (anonymous):

A 12inch pizza cost $8. How much would you charge for a 9inch pizza if you wanted to charge the same amount per bite for both pizzas?

OpenStudy (akashdeepdeb):

12inch pizza = 8 [I love pizza] 9 inch pizza = ??? So as pizza inch increases, amount increases and vice versa. So they vary DIRECTLY. 12/8 = 9/x [Letting x be price for 9 inch pizza] Getting this? :)

OpenStudy (anonymous):

i did it & got 6 before i even posted this & my online class says that it is wrong, thats why im so confused.

OpenStudy (anonymous):

Well, the number if bites depends on the area. A pizza is a circle so the area is \[ A=\pi r^2 \]Where \(r\) is the radius. If you wanted to use diameter \(d\) you'd have \[ A = \pi \left(\frac d2\right)^2=\frac \pi 4 d^2 \]

OpenStudy (anonymous):

Wanting price \(P\) to be proportional that would mean \[ \frac{P_2}{P_1}=\frac{A_2}{A_2}=\frac{(d_2)^2}{(d_1)^2} \]So try solving for \(x\) where: \[ \frac{x}{8}=\frac{9^2}{12^2} \]

OpenStudy (jdoe0001):

is it a squared pizza or a circular one? if I gather it's a circular one, I'd assume 12 stands for the diameter, thus the Area of the 1st one will be \(\bf 2\pi \cdot 6\) the Area of the 2nd one with a diameter of \(\bf 2\pi \cdot 4.5\)

OpenStudy (anonymous):

It doesn't actually matter if it is squared or circular since the relationship still varies directly with the square of the diameter/radius/side etc.

OpenStudy (jdoe0001):

so.... we can say that \(\begin{array}{llll} Area & price\\ 12\pi& 8\\ 9\pi & x \end{array}\bf \implies \cfrac{12\pi}{9\pi} = \left(\cfrac{8}{x}\right)^2\implies \cfrac{12\pi}{9\pi} =\cfrac{8^2}{x^2}\)